# Polychromic Objectives for Reinforcement Learning

### v1

@@ Abstract
\$*\$\$*\$footnotetext: Equal contribution. Correspondence to {jubayer, ifdi1101}@stanford.edu.
@@ 1 Introduction
Reinforcement learning fine-tuning (RLFT) is widely used to enhance the performance of pretrained models across diverse downstream domains. For instance, RLFT has been applied to steer large language models (LLMs) toward instruction following and complex reasoning [Ouyang et al., 2022; DeepSeek-AI et al., 2025; OpenAI et al., 2024; Qwen, 2025]. A common thread across these settings is the availability of expressive generative models (i.e., pretrained distributions), trained on large and diverse datasets, that already exhibit a broad repertoire of strategies. RLFT then refines these distributions by reinforcing the strategies that yield higher reliability and performance.
However, exploration during RLFT remains a central challenge. Prior work [Cui et al., 2025; Zhao et al., 2025] has documented entropy collapse: instead of expanding their repertoire, fine-tuned policies concentrate probability mass on a narrow set of high-reward behaviors already present in the pretrained distribution, effectively sacrificing entropy and diversity. This limits exploration and prevents the discovery of alternative strategies that could expand the base model’s capabilities. Empirically, this effect is captured by the pass@$k$ metric, which measures the probability that at least one out of $k$ independently sampled rollouts succeeds. When $k$ is large, RL-fine-tuned models often underperform their pretrained counterparts because the latter retain greater diversity [Yue et al., 2025; Wu et al., 2025]. Such diversity is practically important; it supports generalization to new tasks [Kumar et al., 2020] and amplifies test-time compute scaling [Snell et al., 2024].
@@
To this end, we propose set reinforcement learning, where the objective is defined over a set of trajectories sampled independently and evaluated by a multi-sample objective [Tang et al., 2025]. Unlike standard RL, which maximizes the likelihood of a single optimal trajectory, set RL maximizes the likelihood of an optimal set of trajectories under a set-level objective. Within this framework, we introduce polychromic objectives which combine reward and diversity by scoring sets highly only if they contain both successful and diverse trajectories. Optimizing a policy with respect to this objective is a principled approach towards encouraging the policy to explore and search for a diverse set of generations that also maximize reward. We then instantiate one such objective and show how proximal policy optimization (PPO) [Schulman et al., 2017b] can be adapted to optimize it effectively, yielding a practical algorithm we call polychromic PPO. We evaluate our method on BabyAI [Chevalier-Boisvert et al., 2019], Minigrid [Chevalier-Boisvert et al., 2023], and Algorithmic Creativity [Nagarajan et al., 2025]. Our results show that polychromic PPO achieves higher rewards and success rates, generates diverse trajectories that substantially improve pass@$k$ coverage, and generalizes more robustly to perturbations in the initial state.
@@ 2 Preliminaries
We consider a Markov decision process defined by state space $\mathcal{S}$, action space $\mathcal{A}$, transition dynamics distribution $p(s_{t+1}\mid s_{t},a_{t})$, reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, initial state distribution $\rho_{0}$ and discount factor $\gamma\in(0,1)$. In reinforcement learning (RL), the goal is to learn a policy that maximizes the value $V(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[R(\tau)\right]$ where $R(\tau)$ is the (discounted) sum of rewards in trajectory $\tau$. The following, known as the performance difference lemma, is a useful result [Kakade & Langford, 2002]:
$V_{\pi_{\theta}}(s_{0})-V_{\pi_{\beta}}(s_{0})=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi}(\cdot\mid s_{0}),a\sim\pi(\cdot\mid s)}\left[A^{\pi_{\beta}}(s,a)\right].$ (1)
This shows that any update from policy $\pi_{\beta}$ to policy $\pi_{\theta}$ such that, at all states visited by $\pi_{\theta}$, the actions taken by $\pi_{\theta}$ yield positive advantage $A^{\pi_{\beta}}(s,a)>0$ will ensure that $\pi_{\theta}$ achieves strictly better performance i.e. $V({\pi_{\theta}})>V({\pi_{\beta}})$. One widely used RL algorithm is proximal policy optimization (PPO) [Schulman et al., 2017b] which, iteratively, collects rollouts under a behavior policy $\pi_{\beta}$ and updates a policy $\pi_{\theta}$ by constraining the divergence between the two policies from growing too large: letting $r_{t}={\pi_{\theta}(a_{t}\mid s_{t})}/{\pi_{\beta}(a_{t}\mid s_{t})}$, PPO optimizes
@@
where $d^{\pi_{\beta}}(s)$ is the stationary state-visitation distribution under policy $\pi_{\beta}$. Here, we use importance sampling to use actions sampled from $\pi_{\beta}$. To use states from the visitation distribution of $\pi_{\beta}$, we clip the ratios to keep the divergence small [Schulman et al., 2017a].
@@ 3.1 Set Reinforcement Learning
This setup contrasts with Tang et al. [2025], which employs trajectory-specific baselines leading to leave-one-out advantages of the form $f(s_{0},\tau_{1:n})-f(s_{0},\tau_{1:i-1},\tau_{i+1:n})$; the update for trajectory $\tau_{i}$ depends on a baseline computed from the remaining trajectories, yielding individualized credit assignment. In our case, a uniform baseline provides a common update signal to all trajectories in the set, enabling optimization with respect to the quality of the entire set of trajectories. In other words, by definition, set RL does not distinguish between trajectories within a set, but instead optimizes the policy by comparing across sets as a whole. Note that the set RL framework can still be used to optimize the standard RL objective by choosing $f(s_{0},\tau_{1:n})=\frac{1}{n}\sum_{i=1}^{n}R(\tau_{i})$. However, the framework allows for a broader class of objectives, such as inference-time objectives  [Tang et al., 2025] and objectives that induce exploration as we will show in §3.2.
@@
These definitions help us better understand the objective of set reinforcement learning. In standard reinforcement learning, we want to learn the policy such that the expected return from a trajectory is maximized. In set reinforcement learning, we want to learn the policy that maximizes the expected return from a tree generated using our policy is maximized. Given these definitions, we have the following result which is an extension of the performance difference lemma  [Kakade & Langford, 2002] to the set reinforcement learning setting (see section C.1 for proof):
@@ 3.2 A Practical Polychromic Objective
Various diversity metrics have been studied and incorporated in reward functions in prior works; examples include the Vendi Score  [Friedman & Dieng, 2023] and classifier-guided diversity  [Zhang et al., 2025; Li et al., 2025]. Our algorithm is designed to be agnostic to the choice of metric: given any diversity function, we evaluate the success and diversity of a set of trajectories and optimize the policy to maximize both in sets.
@@ 3.3 Polychromic PPO
A direct implementation of the definition of set advantage functions as in Lemma 3.2 would require sampling $n$ actions from every visited state, leading to exponential data requirements. To avoid this, we instead rely on vine sampling [Schulman et al., 2017a; Kazemnejad et al., 2025] for on-policy data collection. In vine sampling, after collecting an initial set of rollouts, we select a subset $\{s_{1},\ldots,s_{p}\}$ of the states visited, called rollout states. At each rollout state $s_{i}$, we generate $N$ additional rollouts starting from $s_{i}$. This procedure ensures that we obtain multiple states with independently sampled trajectory sets stemming out of them. The particular scheme we use is closely related to vine TRPO [Schulman et al., 2017a]; details are deferred to §A.1.1. Our algorithm, however, is compatible with any vine-sampling method that guarantees sufficient vine coverage. Note that, since vine sampling requires the ability to reset the environment, our algorithm is only applicable to environments where such resets are possible.
@@
For non-rollout states, the update remains the same as standard PPO, using generalized advantage estimation (GAE) [Schulman et al., 2018]. As is often used in practical implementations of PPO, we also include a per-state KL penalty $D_{\mathrm{KL}}\big(\pi_{\beta}(\cdot\mid s)|\pi_{\theta}(\cdot\mid s)\big)$ at every state visited, which we found helpful for stability. The pseudocode is presented in algorithm 1, with modifications relative to PPO highlighted; extended pseudocode and implementation details are given in appendix A. We report all hyperparameters in §A.1.2. Note that the size of sets $n$ used in set RL and the number of trajectories $N$ generated from rollout states are fixed hyperparameters (in all our experiments, we used $n=4$ and $N=8$).
@@ 4 Experimental Evaluation
Our experiments aim to answer two questions: (1) How does polychromic PPO, a set RL algorithm that explicitly encourages diverse trajectory generation, affect performance? More specifically, does improving diversity come at a significant cost in accuracy and success rate? (2) Does polychromic PPO encourage the policy to explore and learn diverse behaviors? In particular, learning to solve a task through diverse generations should, ideally, increase the pass@$k$ performance i.e., the probability of succeeding at least once when given multiple attempts. (3) Does encouraging the policy to explore and maximize the diversity of generated trajectories help the policy to be more robust to perturbations in the state-visitation distribution? To address these questions, we evaluate on Minigrid [Chevalier-Boisvert et al., 2023], BabyAI [Chevalier-Boisvert et al., 2019], and Algorithmic Creativity [Nagarajan et al., 2025].
@@
Minigrid and BabyAI are grid-world platforms with multiple rooms connected by locked doors and populated with keys, balls, and distractors. Agents receive natural language goals ranging from simple (for example, go to the red ball) to highly compositional (for example, open a red door and then go to the ball on your left after placing the grey ball next to a door). We use a language-conditioned convolutional neural network (CNN) policy that is pretrained on expert demonstrations provided in the official datasets  [Chevalier-Boisvert et al., 2023, 2019]. We then fine-tune and evaluate on 50 fixed configurations (each configuration specifies a grid layout and mission pair).
@@
Algorithmic Creativity is a set of tasks for quantifying model creativity  [Nagarajan et al., 2025]. In the triangle discovery task, an agent must autoregressively output sequences of triangles from undirected graphs that contain many triangles. Solving the task in each graph requires recall and combinations of past knowledge, as the graph is not revealed in-context, but learned during pretraining and through interactions in RLFT. We pretrain on data drawn from 10 graphs and fine-tune on 3 graphs. Here, the verifiable reward is whether the generated sequence is a valid triangle in the graph.
@@
We compare polychromic PPO (Poly-PPO) with REINFORCE with baseline [Williams, 1992] and standard PPO [Schulman et al., 2017b]. Furthermore, we compare with a UCB-style regularization  [Azar et al., 2017] where we add $\lambda_{\mathrm{UCB}}\cdot\min\{1,N(s,a)^{-\frac{1}{2}}\}$ to every advantage $\hat{A}(s,a)$. Here, $N(s,a)$ is the number of times that action $a$ was sampled from state $s$ and $\lambda_{\mathrm{UCB}}$ is a hyperparameter we tune for each method. For polychromic PPO, we define the diversity function $d(s,\tau_{1:n})$ to be the fraction of semantically distinct trajectories in $\tau_{1:n}$; in Minigrid/BabyAI, two trajectories are called distinct if they visit different sets of rooms and, in Algorithmic Creativity, two trajectories are distinct if they visit different sets of nodes. In both cases, $d=0$ if all trajectories visit the same set of rooms or nodes. More details on the environment and our implementation can be found in §A. Throughout all settings, we generate $N=8$ trajectories at rollouts states for vine sampling and use sets of size $n=4$ for poly-PPO (see algorithm 1).
@@ 4.1 How Does Polychromic PPO Affect Performance?
Table 1: Average reward and success rate (%) on BabyAI tasks. Each value is averaged over 100 rollouts across 50 configurations and 3 random seeds.
@@
Results for Algorithmic Creativity are summarized in fig. 2. We rollout 100 times across the 3 graphs seen during RLFT. The creativity metric is defined as the percentage of generations that are unique, valid triangles that were not seen in the pretraining data [Nagarajan et al., 2025]. We also report validity (number of valid triangles constructed), and diversity (unique number of valid triangles). PPO substantially increase validity compared to the pretrained policy, but lower creativity and diversity. On the other hand, polychromic PPO achieves slightly lower validity than standard PPO, but the gap is modest and highlights the trade-off it strikes between success and exploration which we discuss in the next subsection.
4.2 Does Polychromic PPO Encourage Diverse Generations?
@@ 4.3 Does polychromic PPO generalize to state perturbations?
Table 2: Average pass rate (%) in one attempt on BabyAI tasks under large initial-state perturbations.
@@ 5 Entropy Analysis
In this section, we analyze how the entropy of a policy evolves when trained to optimize the polychromic objective in eq. 7. Our guiding question is: Under the polychromic objective, on which actions is a policy most likely to collapse its probability mass? In our analysis, we restrict our attention to the setting with time horizon $H=1$, with binary rewards. We assume a discrete action space and that the diversity function $d(s,a_{1:n})$ equals the fraction of actions in $a_{1:n}$ that are distinct ($d=0$ if the set is a singleton). We assume that our policy has a softmax parameterization. Before turning to the polychromic objective itself, we extend the entropy analysis of Cui et al. [2025] to the set RL setting. Given any set objective $f:\mathcal{S}\times\mathcal{A}^{n}\rightarrow\mathbb{R}$, we ask: how does the entropy of the policy change after one-step update (from $\pi_{\theta}^{k}$ to $\pi_{\theta}^{k+1}$) when learning under this set-RL framework? The following proposition characterizes the first-order change:
@@
The scaffold represents, for every action set $a_{1:n}$, a measure of the policy’s propensity to collapse its entropy around that set. With this apparatus in hand, we next show that the polychromic objective rules out collapse onto homogeneous sets of actions (proof in §C.3):

### v2

@@ 1 Introduction
Reinforcement learning fine-tuning (RLFT) is widely used to enhance the performance of pretrained models across diverse downstream domains. For instance, RLFT has been applied to steer large language models (LLMs) toward instruction following and complex reasoning [23, 6, 22, 25]. A common thread across these settings is the availability of expressive generative models (i.e., pretrained distributions), trained on large and diverse datasets, that already exhibit a broad repertoire of strategies. RLFT then refines these distributions by reinforcing the strategies that yield higher reliability and performance.
However, exploration during RLFT remains a central challenge. Prior work [5, 44] has documented entropy collapse: instead of expanding their repertoire, fine-tuned policies concentrate probability mass on a narrow set of high-reward behaviors already present in the pretrained distribution, effectively sacrificing entropy and diversity. This limits exploration and prevents the discovery of alternative strategies that could expand the base model’s capabilities. Empirically, this effect is captured by the pass@$k$ metric, which measures the probability that at least one out of $k$ independently sampled rollouts succeeds. When $k$ is large, RL-fine-tuned models often underperform their pretrained counterparts because the latter retain greater diversity [42, 39]. Such diversity is practically important; it supports generalization to new tasks [16] and amplifies test-time compute scaling [31].
@@
To this end, we propose set reinforcement learning, where the objective is defined over a set of trajectories sampled independently and evaluated by a multi-sample objective [36]. Unlike standard RL, which maximizes the likelihood of a single optimal trajectory, set RL maximizes the likelihood of an optimal set of trajectories under a set-level objective. Within this framework, we introduce polychromic objectives which combine reward and diversity by scoring sets highly only if they contain both successful and diverse trajectories. Optimizing a policy with respect to this objective is a principled approach towards encouraging the policy to explore and search for a diverse set of generations that also maximize reward. We then instantiate one such objective and show how proximal policy optimization (PPO) [29] can be adapted to optimize it effectively, yielding a practical algorithm we call polychromic PPO. We evaluate our method on BabyAI [3], Minigrid [3], and Algorithmic Creativity [21]. Our results show that polychromic PPO achieves higher rewards and success rates, generates diverse trajectories that substantially improve pass@$k$ coverage, and generalizes more robustly to perturbations in the initial state.
@@ 2 Preliminaries
We consider a Markov decision process defined by state space $\mathcal{S}$, action space $\mathcal{A}$, transition dynamics distribution $p(s_{t+1}\mid s_{t},a_{t})$, reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, initial state distribution $\rho_{0}$ and discount factor $\gamma\in(0,1)$. In reinforcement learning (RL), the goal is to learn a policy that maximizes the value $V(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[R(\tau)\right]$ where $R(\tau)$ is the (discounted) sum of rewards in trajectory $\tau$. The following, known as the performance difference lemma, is a useful result [13]:
$V_{\pi_{\theta}}(s_{0})-V_{\pi_{\beta}}(s_{0})=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi_{\theta}}(\cdot\mid s_{0}),a\sim\pi_{\theta}(\cdot\mid s)}\left[A^{\pi_{\beta}}(s,a)\right].$ (1)
This shows that any update from policy $\pi_{\beta}$ to policy $\pi_{\theta}$ such that, at all states visited by $\pi_{\theta}$, the actions taken by $\pi_{\theta}$ yield positive advantage $A^{\pi_{\beta}}(s,a)>0$ will ensure that $\pi_{\theta}$ achieves strictly better performance i.e. $V({\pi_{\theta}})>V({\pi_{\beta}})$. One widely used RL algorithm is proximal policy optimization (PPO) [29] which, iteratively, collects rollouts under a behavior policy $\pi_{\beta}$ and updates a policy $\pi_{\theta}$ by constraining the divergence between the two policies from growing too large: letting $r_{t}={\pi_{\theta}(a_{t}\mid s_{t})}/{\pi_{\beta}(a_{t}\mid s_{t})}$, PPO optimizes
@@
where $d^{\pi_{\beta}}(s)$ is the stationary state-visitation distribution under policy $\pi_{\beta}$. Here, we use importance sampling to use actions sampled from $\pi_{\beta}$. To use states from the visitation distribution of $\pi_{\beta}$, we clip the ratios to keep the divergence small [27].
@@ 3.1 Set Reinforcement Learning
This setup contrasts with [36], which employs trajectory-specific baselines leading to leave-one-out advantages of the form $f(s_{0},\tau_{1:n})-f(s_{0},\tau_{1:i-1},\tau_{i+1:n})$; the update for trajectory $\tau_{i}$ depends on a baseline computed from the remaining trajectories, yielding individualized credit assignment. In our case, a uniform baseline provides a common update signal to all trajectories in the set, enabling optimization with respect to the quality of the entire set of trajectories. In other words, by definition, set RL does not distinguish between trajectories within a set, but instead optimizes the policy by comparing across sets as a whole. Note that the set RL framework can still be used to optimize the standard RL objective by choosing $f(s_{0},\tau_{1:n})=\frac{1}{n}\sum_{i=1}^{n}R(\tau_{i})$. However, the framework allows for a broader class of objectives, such as inference-time objectives  [36] and objectives that induce exploration as we will show in §3.2.
@@
These definitions help us better understand the objective of set reinforcement learning. In standard reinforcement learning, we want to learn the policy such that the expected return from a trajectory is maximized. In set reinforcement learning, we want to learn the policy that maximizes the expected return from a tree generated using our policy is maximized. Given these definitions, we have the following result which is an extension of the performance difference lemma  [13] to the set reinforcement learning setting (see Section C.1 for proof):
@@ 3.2 A Practical Polychromic Objective
Various diversity metrics have been studied and incorporated in reward functions in prior works; examples include the Vendi Score  [8] and classifier-guided diversity  [43, 18]. Our algorithm is designed to be agnostic to the choice of metric: given any diversity function, we evaluate the success and diversity of a set of trajectories and optimize the policy to maximize both in sets.
@@ 3.3 Polychromic PPO
A direct implementation of the definition of set advantage functions as in Lemma 3.2 would require sampling $n$ actions from every visited state, leading to exponential data requirements. To avoid this, we instead rely on vine sampling [27, 15] for on-policy data collection. In vine sampling, after collecting an initial set of rollouts, we select a subset $\{s_{1},\ldots,s_{p}\}$ of the states visited, called rollout states. At each rollout state $s_{i}$, we generate $N$ additional rollouts starting from $s_{i}$. This procedure ensures that we obtain multiple states with independently sampled trajectory sets stemming out of them. The particular scheme we use is closely related to vine TRPO [27]; details are deferred to §A.1.1. Our algorithm, however, is compatible with any vine-sampling method that guarantees sufficient vine coverage. Note that, since vine sampling requires the ability to reset the environment, our algorithm is only applicable to environments where such resets are possible.
@@
For non-rollout states, the update remains the same as standard PPO, using generalized advantage estimation (GAE) [28]. As is often used in practical implementations of PPO, we also include a per-state KL penalty $D_{\mathrm{KL}}\big(\pi_{\beta}(\cdot\mid s)|\pi_{\theta}(\cdot\mid s)\big)$ at every state visited, which we found helpful for stability. The pseudocode is presented in Algorithm 1, with modifications relative to PPO highlighted; extended pseudocode and implementation details are given in Appendix A. We report all hyperparameters in §A.1.2. Note that the size of sets $n$ used in set RL and the number of trajectories $N$ generated from rollout states are fixed hyperparameters (in all our experiments, we used $n=4$ and $N=8$).
Note that vine sampling is not a necessary component of the algorithm. Consider an environment with a fixed initial state. In this case, if we sample rollouts independently without any vine sampling, we will naturally get multiple rollouts from the fixed initial state allowing us to apply the polychromic advantage to the initial state. Vine sampling simply enables us to apply the polychromic advantage to a larger number of states (i.e., all the rollout states) which helps encourage exploration at various stages of a rollout.
@@ 4 Experimental Evaluation
Our experiments aim to answer two questions: (1) How does polychromic PPO, a set RL algorithm that explicitly encourages diverse trajectory generation, affect performance? More specifically, does improving diversity come at a significant cost in accuracy and success rate? (2) Does polychromic PPO encourage the policy to explore and learn diverse behaviors? In particular, learning to solve a task through diverse generations should, ideally, increase the pass@$k$ performance i.e., the probability of succeeding at least once when given multiple attempts. (3) Does encouraging the policy to explore and maximize the diversity of generated trajectories help the policy to be more robust to perturbations in the state-visitation distribution? To address these questions, we evaluate on Minigrid [3], BabyAI [3], and Algorithmic Creativity [21].
@@
Minigrid and BabyAI are grid-world platforms with multiple rooms connected by locked doors and populated with keys, balls, and distractors. Agents receive natural language goals ranging from simple (for example, go to the red ball) to highly compositional (for example, open a red door and then go to the ball on your left after placing the grey ball next to a door). We use a language-conditioned convolutional neural network (CNN) policy that is pretrained on expert demonstrations provided in the official datasets  [3]. We then fine-tune and evaluate on 50 fixed configurations (each configuration specifies a grid layout and mission pair).
@@
Algorithmic Creativity is a set of tasks for quantifying model creativity  [21]. In the triangle discovery task, an agent must autoregressively output sequences of triangles from undirected graphs that contain many triangles. Solving the task in each graph requires recall and combinations of past knowledge, as the graph is not revealed in-context, but learned during pretraining and through interactions in RLFT. We pretrain on data drawn from 10 graphs and fine-tune on 3 graphs. Here, the verifiable reward is whether the generated sequence is a valid triangle in the graph.
@@
We compare polychromic PPO (Poly-PPO) with REINFORCE with baseline [38] and standard PPO [29]. Furthermore, we compare with a UCB-style regularization  [1] where we add $\lambda_{\mathrm{UCB}}\cdot\min\{1,N(s,a)^{-\frac{1}{2}}\}$ to every advantage $\hat{A}(s,a)$. Here, $N(s,a)$ is the number of times that action $a$ was sampled from state $s$ and $\lambda_{\mathrm{UCB}}$ is a hyperparameter we tune for each method. For polychromic PPO, we define the diversity function $d(s,\tau_{1:n})$ to be the fraction of semantically distinct trajectories in $\tau_{1:n}$; in Minigrid/BabyAI, two trajectories are called distinct if they visit different sets of rooms and, in Algorithmic Creativity, two trajectories are distinct if they visit different sets of nodes. In both cases, $d=0$ if all trajectories visit the same set of rooms or nodes. More details on the environment and our implementation can be found in §A. Throughout all settings, we generate $N=8$ trajectories at rollouts states for vine sampling and use sets of size $n=4$ for poly-PPO (see Algorithm 1).
@@ 4.1 How Does Polychromic PPO Affect Performance?
Table 1: Performance on BabyAI and MiniGrid tasks. Each entry is reported as $(x,y)$, where $x$ is the average episodic reward and $y$ is the success rate (in %). Results are averaged over 100 rollouts across 50 configurations and 3 random seeds.
@@
Results for Algorithmic Creativity are summarized in Fig. 2. We rollout 100 times across the 3 graphs seen during RLFT. The creativity metric is defined as the percentage of generations that are unique, valid triangles that were not seen in the pretraining data [21]. We also report validity (number of valid triangles constructed), and diversity (unique number of valid triangles). PPO substantially increase validity compared to the pretrained policy, but lower creativity and diversity. On the other hand, polychromic PPO achieves slightly lower validity than standard PPO, but the gap is modest and highlights the trade-off it strikes between success and exploration which we discuss in the next subsection.
4.2 Does Polychromic PPO Encourage Diverse Trajectories?
@@ 4.3 Does polychromic PPO generalize to state perturbations?
Table 2: Average pass rate (%) in one attempt on BabyAI and Minigrid tasks under large initial-state perturbations.
@@ 5 Entropy Analysis
In this section, we analyze how the entropy of a policy evolves when trained to optimize the polychromic objective in Eq. 7. Our guiding question is: Under the polychromic objective, on which actions is a policy most likely to collapse its probability mass? In our analysis, we restrict our attention to the setting with time horizon $H=1$, with binary rewards. We assume a discrete action space and that the diversity function $d(s,a_{1:n})$ equals the fraction of actions in $a_{1:n}$ that are distinct ($d=0$ if the set is a singleton). We assume that our policy has a softmax parameterization. Before turning to the polychromic objective itself, we extend the entropy analysis of [5] to the set RL setting. Given any set objective $f:\mathcal{S}\times\mathcal{A}^{n}\rightarrow\mathbb{R}$, we ask: how does the entropy of the policy change after one-step update (from $\pi_{\theta}^{k}$ to $\pi_{\theta}^{k+1}$) when learning under this set-RL framework? The following proposition characterizes the first-order change:
@@
The scaffold represents, for every action set $a_{1:n}$, a measure of the policy’s propensity to collapse its entropy around that set. We illustrate this further through the following lemma which shows us how the scaffold value of an action set affects the change in the probability of a policy sampling the set (proof in §C.3):
Lemma 5.3
Consider any set of actions $a_{1:n}\in\mathcal{A}^{n}$. The change in the log probability of sampling this set of actions after one policy update using set RL can be written as the following first-order approximation:
$\log\pi_{\theta}^{k+1}(a_{1:n}\mid s)\approx\log\pi_{\theta}^{k}(a_{1:n}\mid s)+\lambda\Lambda_{f}(a_{1:n};\pi_{\theta}^{k})-\lambda C(\theta^{k})$
where $C(\theta^{k})$ is a function independent of $a_{1:n}$, and $\lambda$ and $\lambda^{\prime}$ are constants.
With this apparatus in hand, we next show that the polychromic objective rules out collapse onto homogeneous sets of actions (proof in §C.4):